OpenStudy (anonymous):

A 0.5 kg ball moves in a circle that is 0.4 m in radius at a speed of 4.0 m/s. Calculate its centripetal acceleration. (40 m/s2) i dont undertstand how to set up the problem?

OpenStudy (jfraser):

do you have an equation that defines centripetal acceleration?

OpenStudy (anonymous):

yeah hold on

OpenStudy (anonymous):

here are my equations The formula for calculating centripetal acceleration is ac = v2 / r v = velocity r = radius of circle Distance around = Circumference C = 2πr T = time for one full cycle (period) v = 2πr / T

OpenStudy (jfraser):

so you've got a velocity and a radius. Look at the first equation

OpenStudy (anonymous):

so i would just divide v2 by the raduis then?

OpenStudy (anonymous):

so 4.0(4.0)/0.4)=16/0.4=40?

OpenStudy (mrnood):

it's good to put the units in with a quantity - it helps to check your calculations. What are the units of acceleration?

OpenStudy (anonymous):

the units are m/s2

OpenStudy (mrnood):

so the full answer is 40m/s^2 :-)

OpenStudy (anonymous):

thanks you:)